OpenStudy (anonymous):

Simplify in Standard Form: 192=(x+1)(x-4)(x) PLEASE help thank you So much

OpenStudy (anonymous):

standard for = ax+by=c

OpenStudy (anonymous):

i have a question would it be like this @larryboxaplenty

OpenStudy (anonymous):

where did you get the V from? The foil is correct.

OpenStudy (anonymous):

Oh cuz 192 is the volume just ignore that

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

sorry, I don't know.

OpenStudy (anonymous):

its okay so did i do it right?

OpenStudy (mathmale):

larryboxaplenty typed the following: "standard for = ax+by=c," but probably meant:\[y=ax^2 + bx + c\]since the expression you typed in produces a quadratic function when expanded and simplified.

OpenStudy (mathmale):

@yakima+_Love : Let's start over: Simplify in Standard Form: 192=(x+1)(x-4)(x) 1. Use the FOIL method to multiply out (x-1)(x-4). 2. Multiply each term of the result by x. 3. Move the 192 to the right side of the equation by subtracting 192 from both sides of the equation. 4. Rearrange your terms in descending power of x.